# A First Look at Hydrogen Generation in an Ultramafic Rock with Micro-CT and SEM-BEX

### v1

@@ Abstract
Natural hydrogen (H2) generated by water–rock interaction in ultramafic rocks is increasingly recognised as a potentially important primary energy resource, but the pore-scale processes that control the initiation and early transport of a free gas phase remain poorly constrained. Here we present an in situ X-ray micro-tomography experiment in which an ultramafic granular pack of dunnite from West Papua, Indonesia, saturated with KI-doped brine, is heated to 100C with a pore pressure of 4 bar under 10 bar confining pressure inside a micro-CT scanner. Time-resolved 4D imaging captures the transition from a fully liquid-saturated pore space to the appearance and growth of a distinct gas phase after an $\sim$8 h induction period. Bubbles first nucleate near the top of the sample before becoming distributed throughout the imaged volume as a connected ganglia. The nucleating gas phase is most plausibly dominated by molecular hydrogen (H2) generated by low-temperature fluid–rock reaction, as indicated by independent hydrogen-presence detectors, although we cannot yet fully exclude minor contributions from other gases. SEM-BEX imaging reveals textural alteration and local changes in elemental signals between reacted and unreacted material. Taken together, these observations provide spatially and temporally resolved evidence for gas generation during low-temperature alteration of ultramafic grains and demonstrate that pore-scale imaging can directly link water–rock reaction kinetics, gas generation and multiphase flow behaviour in natural hydrogen systems.
@@ 1 Introduction
Natural geologic hydrogen (H2) is emerging as a potentially vast, carbon-free primary energy resource that could complement, or in some settings substitute, engineered “green” H2 production. Global compilations of gas analyses from seeps, boreholes and mines show that H2 is far more widespread than previously appreciated and occurs in many tectonic settings, often as a major component of the gas phase (Zgonnik, 2020). Recent system-scale syntheses formalise the concept of a “H2 system”, in which H2 is generated, migrates and accumulates in traps in an analogous way to hydrocarbons (Jackson and Vinchon, 2024; Prinzhofer and Deville, 2015), and multi-scale reviews highlight the key migration pathways and timescales from the planetary to basin scale (Lodhia et al., 2024). This has led to growing industrial interest in “native” or “natural” hydrogen, with first-order assessments suggesting that alteration-driven systems (e.g. areas of serpentinisation) could sustain commercially significant fluxes over geological timescales (Gaucher, 2020; Evans et al., 2018), and to proposals for engineered “rock-based” hydrogen concepts (white and “orange” hydrogen) that explicitly exploit water–rock reactions in mafic and ultramafic rocks as part of the low-carbon hydrogen portfolio (Osselin et al., 2022).
Among the diverse proposed sources of natural H2, low- to high-temperature serpentinization of ultramafic rocks is generally considered the most efficient and volumetrically important (Oze and Sharma, 2007; McCollom and Bach, 2009; Jackson and Vinchon, 2024). In this process, Fe2+-bearing primary minerals such as olivine and pyroxene react with water to form Fe3+-bearing phases (e.g., magnetite, Fe(III)-serpentine), generating H2 as a by-product (Sleep et al., 2004; Holm et al., 2015). Thermodynamic and kinetic models predict that H2 yields depend sensitively on temperature, water-to-rock ratio, mineralogy, water chemistry, and redox state (McCollom and Bach, 2009; McCollom, 2016; Ely et al., 2023), while field observations from ophiolites and oceanic and continental settings (e.g., Samail ophiolite, Oman; Lost City hydrothermal field) demonstrate that serpentinization can sustain long-lived fluxes of H2-rich, hyperalkaline fluids (Schrenk et al., 2013; Klein et al., 2019). More recently, work in the Samail ophiolite Oman has shown that significant H2 can also be generated during late, low-temperature ($200^{\circ}$C) alteration of already serpentinized peridotite, extending the window of H2 production to cooler, near-surface conditions (Ellison et al., 2021).
@@
Core-scale reactive transport experiments bridge part of this gap by preserving rock fabric and permeability while allowing continuous geochemical monitoring. Percolation of reactive fluids through ultramafic rocks and serpentinites has demonstrated strong couplings between serpentinization, carbonation, permeability reduction and H2 production (e.g., Osselin et al., 2021; Albers et al., 2021; Farough et al., 2016). Yet in most such studies the rock behaves as a “black box”: H2 production is inferred from outlet fluid chemistry or headspace analyses, and any gas phase that forms inside the pore network remains invisible. Even in well-characterised natural systems such as the Samail ophiolite, constraints on H2 generation and transport rely primarily on fluid chemistry, isotopes and mineralogical proxies rather than direct imaging of the gas phase itself (Miller et al., 2017; Ellison et al., 2021). As a result, the fundamental pore-scale mechanisms governing the initiation and early transport of H2, bubble nucleation sites, snap-off events, critical gas saturation and connectivity, are still poorly constrained.
@@
In parallel, a large and rapidly growing body of work has used X-ray micro-tomography (micro-CT) to image dynamic multiphase flow and reactive transport in rocks. Time-resolved (4D) micro-CT has been applied to dissolution and precipitation in carbonates and sandstones, enabling direct observations of evolving pore structures, wormholing and reaction-front dynamics under reservoir conditions (Noiriel and Soulaine, 2021; Noiriel and Renard, 2022; Menke et al., 2015). Similar techniques have been used to quantify residual trapping and ganglion dynamics during CO2–brine flow, to measure capillary pressure–saturation relationships, and to calibrate pore-network and direct numerical simulations (Andrew et al., 2014; Jiang and Tsuji, 2015; Bultreys et al., 2016). These studies demonstrate that 4D imaging can resolve micron-scale fluid–fluid and fluid–solid interfaces, capturing non-linear displacement patterns, snap-off and reconnection events that strongly influence macroscopic transport properties.
@@
Only recently has this pore-scale imaging toolkit been applied to H2 itself. Micro-CT studies of H2–brine flow in sandstones show that H2 behaves as a non-wetting phase in many reservoir-relevant conditions, with significant capillary trapping, strong sensitivity to rock heterogeneity and subtle but important differences from CO2–brine systems (Thaysen et al., 2023; Jangda et al., 2023; Rezaei et al., 2022). These works provide critical benchmarks for underground H2 storage, but they all consider externally injected H2 displacing brine; the H2 is not generated in situ by water–rock reaction. Complementary core-flooding experiments and relative permeability measurements reveal that H2–water flow in rocks is typically capillary-limited, with low minimum water saturations and relative permeability curves that differ from those of CO2–water systems (Yekta et al., 2018; Rezaei et al., 2022; Boon and Hajibeygi, 2022). Reviews of underground H2 storage emphasise that uncertainties in pore-scale trapping, dissolution and reaction with the rock and microbiota are among the key scientific challenges for safe and efficient large-scale deployment (Heinemann et al., 2021; Miocic et al., 2023).
@@ 2.4 Micro-CT Image Analysis
To quantify the evolution of the gas phase, all time-lapse micro-CT volumes were processed using a consistent image-analysis workflow in the image processing software Avizo 2025.1 (Thermo Fisher Scientific). First, each scan was rigidly registered to the initial reference scan by maximising mutual information between grey-scale volumes and resampled onto the reference grid using a Lanczos kernel (Lanczos, 1956) to preserve edge sharpness. Second, we applied a non-local means filter to reduce high-frequency noise while retaining phase boundaries in the greyscale images (Buades et al., 2005). Third, the solid, brine and gas phases were segmented in the Avizo Segmentation Editor using simple global grey-level thresholding, guided by the histogram and visual inspection of representative slices. Conservative thresholds were chosen to separate solid from pore space, and within the pore space voxels with the lowest attenuation were classified as gas. Small isolated gas clusters below a few voxels were removed as likely noise using basic morphological operations. A representative 2D slice illustrating the raw, filtered and segmented images is shown in Figure 2. For each time step we then calculated bulk gas saturation as the ratio of gas-filled voxels to pore-space voxels. Visual inspection of representative slices and 3D renderings was used to check the segmentation quality; we therefore regard the resulting saturation values as approximate but internally consistent across the time series.
@@ 4 Discussion
The delay before the first visible bubbles can be understood as the time needed to build up a sufficient concentration of dissolved H2 in the pore water. From the micro-CT field of view and core-holder geometry, the total sample volume is of order $V_{s}\sim 10^{-7}\,$\mathrm{m}^{3}$\$, and the segmented images give a porosity of $\phi=0.375$, so that the pore volume is $V_{p}\approx\phi V_{s}\approx 4\times 10^{-8}\,$\mathrm{m}^{3}$\$ ($\sim 0.04\,$\mathrm{m}\mathrm{L}$\$). Published measurements and correlations for H2 solubility in pure water and NaCl brines at temperatures near $100^{\circ}$C and H2 partial pressures of order 1–10 bar give equilibrium dissolved concentrations of $c_{\mathrm{sat}}\approx(2$$6)\times 10^{-4}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{L}^{-1}$\$ (e.g., Wiesenburg and Guinasso, 1979; Kling, 1991; Zhu et al., 2022). At these conditions the pore fluid remains entirely in the liquid phase (water vapour pressure at 100C is $\sim$1 bar), so the free gas we observe cannot be attributed to boiling of the brine itself. For our pore volume this corresponds to a total amount of dissolved H2 at near-saturation of
@@
Batch experiments on olivine and partially serpentinized peridotite at similar temperatures report surface-normalised H2 generation rates spanning roughly $10^{-10}$$10^{-8}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{m}^{-2}\,\mathrm{s}^{-1}$\$ depending on mineralogy and fluid composition (e.g., Mayhew et al., 2013; Neubeck et al., 2014; McCollom, 2016). Complementary flow-through experiments on olivine under hydrothermal conditions indicate comparable order-of-magnitude rates while highlighting the influence of fluid flow and evolving permeability on serpentinization and H2 release (Ross et al., 2025). The image-derived grain surface area accessible to the pore fluid within the imaged volume is $A\approx 4.3\times 10^{-4}\,$\mathrm{m}^{2}$\$ (Table 1a), so the implied total H2 production rate is
@@ 4.1 Implications for natural H2 systems
First, the long pre-nucleation period implies that substantial water–rock reaction and H2 production can occur before a free gas phase becomes detectable at the pore scale. In natural settings, zones of active H2 generation may therefore be dominated by dissolved H2 for extended periods, with little or no expression as free gas until local supersaturation and favourable nucleation conditions are met. Experimental work on fractured ultramafic rocks likewise shows that serpentinization can rapidly seal existing flow paths by reducing fracture permeability by orders of magnitude, implying that sustained fluid and gas transport requires continual generation of new permeability (Farough et al., 2016). This behaviour is directly relevant to proposals for exploiting rock-based hydrogen resources, including “orange hydrogen” schemes that couple water–rock–H2 generation with carbon mineralisation and storage (Osselin et al., 2022).
Second, once nucleation thresholds are crossed, the rapid transition to a connected gas ganglia suggests that gas connectivity may be established over relatively short timescales compared to the overall reaction history. This is consistent with the view that small changes in capillary entry pressures or wettability can trigger large changes in effective gas mobility (e.g., Thaysen et al., 2023; Jangda et al., 2023; Rezaei et al., 2022). In heterogeneous natural reservoirs, such threshold behaviour is likely to be highly localised, so migration pathways for natural H2 may be controlled by a limited number of “activated” reaction zones.
Third, the modest changes between 23 and 30 h highlight that, once a gas phase is established, further H2 production can be partitioned between dissolution into the surrounding fluid and incremental growth of existing bubbles rather than continued nucleation. This behaviour is directly relevant to both natural H2 accumulations and underground H2 storage, where the balance between dissolution, trapping and mobile gas strongly influences storage security and recoverable volumes (Heinemann et al., 2021; Miocic et al., 2023). Our results emphasise that predicting this balance requires not only bulk reaction kinetics, but also an understanding of pore-scale nucleation thresholds and the evolving microstructure of reactive rocks.

### v2

@@ Abstract
Natural hydrogen (H2) generated by water–rock interaction in ultramafic rocks is increasingly recognised as a potentially important primary energy resource, but the pore-scale processes that control the initiation and early transport of a free gas phase remain poorly constrained. Here we present an in situ X-ray micro-tomography experiment in which an ultramafic granular pack of ultramafic from Papua Province, Indonesia, saturated with KI-doped brine, is heated to 100C with a pore pressure of 4 bar under 10 bar confining pressure inside a micro-CT scanner. Time-resolved 4D imaging captures the transition from a fully liquid-saturated pore space to the appearance and growth of a distinct gas phase after an $\sim$8 h induction period. Bubbles first nucleate near the top of the sample before becoming distributed throughout the imaged volume as a connected ganglia. The nucleating gas phase is most plausibly dominated by molecular hydrogen (H2) generated by low-temperature fluid–rock reaction, as indicated by independent hydrogen-presence detectors, although we cannot yet fully exclude minor contributions from other gases. SEM-BEX imaging reveals textural alteration and local changes in elemental signals between reacted and unreacted material. Taken together, these observations provide spatially and temporally resolved evidence for gas generation during low-temperature alteration of ultramafic grains and demonstrate that pore-scale imaging can directly link water–rock reaction kinetics, gas generation and multiphase flow behaviour in natural hydrogen systems.
@@ 1 Introduction
Natural geologic hydrogen (H2) is emerging as a potentially vast, carbon-free primary energy resource that could complement, or in some settings substitute, engineered “green” H2 production. Global compilations of gas analyses from seeps, boreholes and mines show that H2 is far more widespread than previously appreciated and occurs in many tectonic settings, often as a major component of the gas phase (Zgonnik2020). Recent system-scale syntheses formalise the concept of a “H2 system”, in which H2 is generated, migrates and accumulates in traps in an analogous way to hydrocarbons (Jackson2024; Prinzhofer2015), and multi-scale reviews highlight the key migration pathways and timescales from the planetary to basin scale (Lodhia2024_H2Migration). This has led to growing industrial interest in “native” or “natural” hydrogen, with first-order assessments suggesting that alteration-driven systems (e.g. areas of serpentinisation) could sustain commercially significant fluxes over geological timescales (Gaucher2020), and to proposals for engineered “rock-based” hydrogen concepts (white and “orange” hydrogen) that explicitly exploit water–rock reactions in mafic and ultramafic rocks as part of the low-carbon hydrogen portfolio (Osselin2022).
Among the diverse proposed sources of natural H2, low- to high-temperature serpentinization of ultramafic rocks is generally considered the most efficient and volumetrically important (Oze2007; McCollom2009; Jackson2024). In this process, Fe2+-bearing primary minerals such as olivine and pyroxene react with water to form Fe3+-bearing phases (e.g., magnetite, Fe(III)-serpentine), generating H2 as a by-product (Sleep2004; Holm2015). Thermodynamic and kinetic models predict that H2 yields depend sensitively on temperature, water-to-rock ratio, mineralogy, water chemistry, and redox state (McCollom2009; McCollom2016; Ely2023), while field observations from ophiolites and oceanic and continental settings (e.g., Samail ophiolite, Oman; Lost City hydrothermal field) demonstrate that serpentinization can sustain long-lived fluxes of H2-rich, hyperalkaline fluids (Schrenk2013). More recently, work in the Samail ophiolite Oman has shown that significant H2 can also be generated during late, low-temperature ($200^{\circ}$C) alteration of already serpentinized peridotite, extending the window of H2 production to cooler, near-surface conditions (Ellison2021).
@@
Core-scale reactive transport experiments bridge part of this gap by preserving rock fabric and permeability while allowing continuous geochemical monitoring. Percolation of reactive fluids through ultramafic rocks and serpentinites has demonstrated strong couplings between serpentinization, carbonation, permeability reduction and H2 production (e.g., Wang2019; Albers2021; Farough2016). Yet in most such studies the rock behaves as a “black box”: H2 production is inferred from outlet fluid chemistry or headspace analyses, and any gas phase that forms inside the pore network remains invisible. Even in well-characterised natural systems such as the Samail ophiolite, constraints on H2 generation and transport rely primarily on fluid chemistry, isotopes and mineralogical proxies rather than direct imaging of the gas phase itself (Miller2016; Ellison2021). As a result, the fundamental pore-scale mechanisms governing the initiation and early transport of H2, bubble nucleation sites, snap-off events, critical gas saturation and connectivity, are still poorly constrained.
@@
In parallel, a large and rapidly growing body of work has used X-ray micro-tomography (micro-CT) to image dynamic multiphase flow and reactive transport in rocks. Time-resolved (4D) micro-CT has been applied to dissolution and precipitation in carbonates and sandstones, enabling direct observations of evolving pore structures, wormholing and reaction-front dynamics under reservoir conditions (NoirielSoulaine2021; NoirielRenard2022; Menke2015). Similar techniques have been used to quantify residual trapping and ganglion dynamics during CO2–brine flow, to measure capillary pressure–saturation relationships, and to calibrate pore-network and direct numerical simulations (Andrew2014; Jiang2015; Bultreys2016). These studies demonstrate that 4D imaging can resolve micron-scale fluid–fluid and fluid–solid interfaces, capturing non-linear displacement patterns, snap-off and reconnection events that strongly influence macroscopic transport properties.
@@
Only recently has this pore-scale imaging toolkit been applied to H2 itself. Micro-CT studies of H2–brine flow in sandstones show that H2 behaves as a non-wetting phase in many reservoir-relevant conditions, with significant capillary trapping, strong sensitivity to rock heterogeneity and subtle but important differences from CO2–brine systems (Thaysen2023; Jangda2023; Rezaei2022). These works provide critical benchmarks for underground H2 storage, but they all consider externally injected H2 displacing brine; the H2 is not generated in situ by water–rock reaction. Complementary core-flooding experiments and relative permeability measurements reveal that H2–water flow in rocks is typically capillary-limited, with low minimum water saturations and relative permeability curves that differ from those of CO2–water systems (Yekta2018; Rezaei2022; Boon2022). Reviews of underground H2 storage emphasise that uncertainties in pore-scale trapping, dissolution and reaction with the rock and microbiota are among the key scientific challenges for safe and efficient large-scale deployment (Heinemann2021; Miocic2023).
@@ 2.4 Micro-CT Image Analysis
To quantify the evolution of the gas phase, all time-lapse micro-CT volumes were processed using a consistent image-analysis workflow in the image processing software Avizo 2025.1 (Thermo Fisher Scientific). First, each scan was rigidly registered to the initial reference scan by maximising mutual information between grey-scale volumes and resampled onto the reference grid using a Lanczos kernel (Lanczos1956) to preserve edge sharpness. Second, we applied a non-local means filter to reduce high-frequency noise while retaining phase boundaries in the greyscale images (Buades2005). Third, the solid, brine and gas phases were segmented in the Avizo Segmentation Editor using simple global grey-level thresholding, guided by the histogram and visual inspection of representative slices. Conservative thresholds were chosen to separate solid from pore space, and within the pore space voxels with the lowest attenuation were classified as gas. Small isolated gas clusters below a few voxels were removed as likely noise using basic morphological operations. A representative 2D slice illustrating the raw, filtered and segmented images is shown in Figure 2. For each time step we then calculated bulk gas saturation as the ratio of gas-filled voxels to pore-space voxels. Visual inspection of representative slices and 3D renderings was used to check the segmentation quality; we therefore regard the resulting saturation values as approximate but internally consistent across the time series.
@@ 4 Discussion
The delay before the first visible bubbles can be understood as the time needed to build up a sufficient concentration of dissolved H2 in the pore water. From the micro-CT field of view and core-holder geometry, the total sample volume is of order $V_{s}\sim 10^{-7}\,$\mathrm{m}^{3}$\$, and the segmented images give a porosity of $\phi=0.375$, so that the pore volume is $V_{p}\approx\phi V_{s}\approx 4\times 10^{-8}\,$\mathrm{m}^{3}$\$ ($\sim 0.04\,$\mathrm{m}\mathrm{L}$\$). Published measurements and correlations for H2 solubility in pure water and NaCl brines at temperatures near $100^{\circ}$C and H2 partial pressures of order 1–10 bar give equilibrium dissolved concentrations of $c_{\mathrm{sat}}\approx(2$$6)\times 10^{-4}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{L}^{-1}$\$ (e.g., Wiesenburg1979_H2Solubility). At these conditions the pore fluid remains entirely in the liquid phase (water vapour pressure at 100C is $\sim$1 bar), so the free gas we observe cannot be attributed to boiling of the brine itself. For our pore volume this corresponds to a total amount of dissolved H2 at near-saturation of
@@
Batch experiments on olivine and partially serpentinized peridotite at similar temperatures report surface-normalised H2 generation rates spanning roughly $10^{-10}$$10^{-8}\,$\mathrm{m}\mathrm{o}\mathrm{l}\,\mathrm{m}^{-2}\,\mathrm{s}^{-1}$\$ depending on mineralogy and fluid composition (e.g., Mayhew2013; Neubeck2014; McCollom2016). Complementary flow-through experiments on olivine under hydrothermal conditions indicate comparable order-of-magnitude rates while highlighting the influence of fluid flow and evolving permeability on serpentinization and H2 release (Ross2025_OlivineFlow). The image-derived grain surface area accessible to the pore fluid within the imaged volume is $A\approx 4.3\times 10^{-4}\,$\mathrm{m}^{2}$\$ (Table 1a), so the implied total H2 production rate is
@@ 4.1 Implications for natural H2 systems
First, the long pre-nucleation period implies that substantial water–rock reaction and H2 production can occur before a free gas phase becomes detectable at the pore scale. In natural settings, zones of active H2 generation may therefore be dominated by dissolved H2 for extended periods, with little or no expression as free gas until local supersaturation and favourable nucleation conditions are met. Experimental work on fractured ultramafic rocks likewise shows that serpentinization can rapidly seal existing flow paths by reducing fracture permeability by orders of magnitude, implying that sustained fluid and gas transport requires continual generation of new permeability (Farough2016). This behaviour is directly relevant to proposals for exploiting rock-based hydrogen resources, including “orange hydrogen” schemes that couple water–rock–H2 generation with carbon mineralisation and storage (Osselin2022).
Second, once nucleation thresholds are crossed, the rapid transition to a connected gas ganglia suggests that gas connectivity may be established over relatively short timescales compared to the overall reaction history. This is consistent with the view that small changes in capillary entry pressures or wettability can trigger large changes in effective gas mobility (e.g., Thaysen2023; Jangda2023; Rezaei2022). In heterogeneous natural reservoirs, such threshold behaviour is likely to be highly localised, so migration pathways for natural H2 may be controlled by a limited number of “activated” reaction zones.
Third, the modest changes between 23 and 30 h highlight that, once a gas phase is established, further H2 production can be partitioned between dissolution into the surrounding fluid and incremental growth of existing bubbles rather than continued nucleation. This behaviour is directly relevant to both natural H2 accumulations and underground H2 storage, where the balance between dissolution, trapping and mobile gas strongly influences storage security and recoverable volumes (Heinemann2021; Miocic2023). Our results emphasise that predicting this balance requires not only bulk reaction kinetics, but also an understanding of pore-scale nucleation thresholds and the evolving microstructure of reactive rocks.
@@ 5 Conclusions
Open Research Section
The micro-CT image volumes, processed segmentations, and representative SEM-BEX images used in this study will be made available in an open-access repository (e.g., Zenodo) upon acceptance of the manuscript. A persistent DOI and metadata will be provided in the final published version.
Conflict of Interest declaration
The authors declare there are no conflicts of interest for this manuscript.